OpenStudy (anonymous):

Evaluate 2 divide x(x + y) if x = 2 and y=4 then how would the work be write with the correct answer

OpenStudy (anonymous):

x = 2 y=4 x(x + y)=2(2+4)=2*6=12 substitute apply----> BODMAS

OpenStudy (anonymous):

thank you correct